Find the de bröglie of a ball of mass 0.12 kg moving with speed 20 m/s

Grade:10

1 Answers

Arun
5 years ago
 m-34/(0.12x20) = 3.412 x 10-34Dear VikashAccording to de Broglie equationWave length = h/mv = 6.255 x 10